OpenStudy (anonymous):

A school district needs 3 teachers for every 70 children. They expect to have 14,700 children in the district next year. At this time, they have 612 teachers. How many more teachers are needed?

OpenStudy (anonymous):

can someone explain how you get 18

jimthompson5910 (jim_thompson5910):

you solve 3/70 = x/14700 for x to get 3/70 = x/14700 3*14700 = 70x 44100 = 70x 70x = 44100 x= 44100/70 x = 630 So 630 teachers are needed. There are currently 612 of them, so you need 630 - 612 = 18 teachers

OpenStudy (anonymous):

70 students=3 teachers or 70s=3t 14700s= (70*210)s (70*210)s=(70s)*210=(3t)*210=630t So, 630 teachers required for 14700 students. There are 612 already present. So, 630-612=18

OpenStudy (anonymous):

for every 70 students, there are 3 teachers So, because there are 14700 students, there needs to be 630 teachers That means that since you already have 612 teachers, you need 18 more